OpenStudy (anonymous):

If a triangle has as an area of 144 ft2 and a height of 24 ft, then what is the base? 1) 6 ft 2) 24 ft 3) 18 ft 4) 12 ft

OpenStudy (anonymous):

144 = 1/2 * 24 * x => x = 12 ft

OpenStudy (anonymous):

Let the base be b. 1/2 * h * b = 144 1/2 * 24 * b = 144 12b = 144 b = 12

OpenStudy (anonymous):

Thanks .

OpenStudy (anonymous):

yw